Known vectors:


a = (α, 0, 1), b = (1, β, 1), c = (1, 1, γ)


Determine the values of α, β,γ when the three vectors are orthogonal to each other.


1
Expert's answer
2021-11-26T12:45:41-0500

"a = (α, 0, 1)=αi+k \\
b = (1, β, 1)=i+βj+k \\
c = (1, 1, γ)=i+j+γk\\"


Since a,b, and care pairwise orthogonal to each other

"\bar{a}*\bar{b}=0 \space \& \space \bar{a}*\bar{c}=0\\
(αi + k)(i+βj+k)=0\\
α+0+1=0\\
α=-1"


"(αi + k)(i+j+γk)=0\\
α+0+γ=0\\
α=-γ\\
γ=1"


"(i + βj+k)(i+j+γk)=0\\
1+β+γ=0\\
1+β+γ=0\\
β=-2"